OpenStudy (anonymous):

Graph the exponential function g(x)=-(3/2)^x . To graph the function, plot the points on the graph with x-values -2,-1 ,0 , 1 , and 2 , draw all asymptotes, and then click on the graph icon.

OpenStudy (pfenn1):

I don't understand what it means by "click on the graph icon" but plug each of the values of x into the equation and solve for g(x). For instance, for x = -2\[g(x)=-\left( \frac32\right)^{-2}=-\left( \frac23\right)^2=-\frac49\]

OpenStudy (anonymous):

im suppost to graph each one of these equations so how do you graph 4/9

OpenStudy (pfenn1):

No, you don't graph each equation, you graph the points. This one is (-2, -4/9)

OpenStudy (anonymous):

now how do i make these point exponential equations?

OpenStudy (anonymous):

This site really helps... https://www.desmos.com/calculator